# ADAM-SINDy: An Efficient Optimization Framework for Parameterized Nonlinear Dynamical System Identification

### v1

@@ 2.2 ADAM Optimization Augmented SINDy: ADAM-SINDy
$\mathcal{L}_{b}\leftarrow\left[\|\dot{\mathbf{X}}_{b}-\mathbf{\Theta}(\mathbf{X}_{b};\mathbf{\Lambda})\thinspace\mathbf{\Xi}\|_{2}^{2}+\lambda\|\mathbf{\Gamma}\thinspace\mathbf{\Xi}\|_{1}\right]$
       Update Parameters:
$\mathbf{\Xi}_{n+1}\leftarrow\mathbf{\Xi}_{n}-\eta_{\mathbf{\Xi}}\thinspace\nabla_{\mathbf{\Xi}}\thinspace\mathcal{L}_{b}(\mathbf{\Theta}_{b},\mathbf{\Xi},\mathbf{\Lambda},\lambda,\mathbf{\Gamma})$
$\mathbf{\Lambda}_{n+1}\leftarrow\mathbf{\Lambda}_{n}-\eta_{\mathbf{\Lambda}}\thinspace\nabla_{\mathbf{\Lambda}}\thinspace\mathcal{L}_{b}(\mathbf{\Theta}_{b},\mathbf{\Xi},\mathbf{\Lambda},\lambda,\mathbf{\Gamma})$
$\lambda_{n+1}\leftarrow\lambda_{n}\thinspace-\eta_{\lambda}\thinspace\nabla_{\lambda}\thinspace\mathcal{L}_{b}(\mathbf{\Theta}_{b},\mathbf{\Xi},\mathbf{\Lambda},\lambda,\mathbf{\Gamma})$
$\mathbf{\Gamma}_{n+1}\leftarrow\mathbf{\Gamma}_{n}\thinspace-\eta_{\mathbf{\Gamma}}\thinspace\nabla_{\mathbf{\Gamma}}\thinspace\mathcal{L}_{b}(\mathbf{\Theta}_{b},\mathbf{\Xi},\mathbf{\Lambda},\lambda,\mathbf{\Gamma})$
       Thresholding: For each element $i$, set $\mathbf{\Xi}_{n+1}^{i},\mathbf{\Lambda}_{n+1}^{i}\leftarrow 0$ if $|\mathbf{\Xi}_{n+1}^{i}|,|\mathbf{\Lambda}_{n+1}^{i}|\leq\epsilon$
@@
 Optimized parameters $\mathbf{\Xi}^{*}$ and $\mathbf{\Lambda}^{*}$
@@ 2.3 Problem Statement
$\mathbf{\Theta}=\begin{bmatrix}1&\mathbf{X}^{\mathcal{A}}&\sin(\mathcal{B}\mathbf{X})&\cos(\mathcal{C}\mathbf{X})&\exp(\mathcal{D}\mathbf{X})&\mathbf{X}\otimes\sin(\mathcal{E}\mathbf{X})&\mathbf{X}\otimes\cos(\mathcal{F}\mathbf{X})&\mathbf{X}\otimes\exp(\mathcal{G}\mathbf{X})\end{bmatrix}\;,$ (8)
where $\mathbf{X}$ denotes the input data, and $\mathcal{A},\mathcal{B},\mathcal{C},\mathcal{D},\mathcal{E},\mathcal{F},\mathcal{G}$ represent the adaptive nonlinear parameters $(\mathbf{\Lambda})$ that are subject to optimization within the ADAM-SINDy framework. Specifically, $\mathcal{A}$ is chosen to be 1, thereby encompassing polynomial terms up to the first order in the master library for all cases. However, in certain numerical experiments, $\mathcal{A}$ is rendered trainable when it corresponds to a non-integer rational exponent. It is noteworthy that the classical SINDy method relies on precise a priori knowledge of these parameters to successfully identify a sparse system.
@@
To investigate these limitations, two variants of the classical SINDy approach are explored: the SINDy-W method and the SINDy-W/o method. The SINDy-W method represents a fortunate scenario (unlikely in practical complex problems) where the precise value of the unknown nonlinear parameter is included within the range of candidate parameters $\mathbf{\Lambda}$. While such a scenario may hold in controlled or well-defined environments, they are often inadequate in real-world engineering applications, where the underlying parameter values may be unknown or difficult to estimate. Conversely, the SINDy-W/o method assumes no prior knowledge of the parametric values and applies a broad range of distributed parameter values to account for a wider spectrum of possibilities. In our benchmark problems, the range of parameters is selected such that the unknown parameter falls within the range. The SINDy-W/o approach is the more realistic practical version of SINDy, however, it often struggles in cases when the predefined parameter ranges fail to contain the “exact value” required to capture the system’s true behavior. This absence of prior knowledge frequently leads to less accurate, overly complex, and non-parsimonious models. Ultimately, these challenges underscore the need for flexibility and adaptability in parameter selection to ensure accurate system identification, motivating the use of the ADAM-SINDy framework, which dynamically adjusts parameter values during optimization and overcomes the limitations of the classical methods’ predefined parameter spaces.
@@
Here, we adopt the approach established in the classical SINDy method regarding the sparsity-promoting parameters $\lambda$ and $\mathbf{\Gamma}$  34. Specifically, $\mathbf{\Gamma}$ is treated as a diagonal identity matrix, with the $\lambda$ parameter being the sole variable used to control sparsity for the two variants considered: classical SINDy-W and SINDy-W/o methods. Notably, the numerical experiments conducted with the ADAM-SINDy framework indicate that the simultaneous optimization and utilization of both sparsity-promoting knobs is not essential; indeed, no significant differences in outcomes were observed when optimizing only one of these knobs.
@@
The ADAM-SINDy methodology is implemented using the PyTorch library, considering both batch-wise and full-batch optimization strategies, and employs the ADAM optimizer to facilitate the simultaneous optimization of all unknown variables involved in the sparse regression problem, including the parameters $\mathbf{\Lambda}$, the coefficients $\mathbf{\Xi}$, and the sparsity knob $\mathbf{\Gamma}$, with the parameter $\lambda$ fixed at a value of 1 and not subject to optimization. In this implementation, the parameter $\mathbf{\Lambda}\in\mathbb{R}^{m\times\hat{p}}$ is defined as a non-square scalar matrix associated with $\hat{p}$ nonlinear candidate parameters corresponding to each of the $m$ state variables. This structure facilitates vectorization, allowing the ADAM optimizer to be invoked only once for all nonlinear parameters rather than for each individual value. Consequently, while $\mathbf{\Lambda}$ parameters behave like scalar values algorithmically, it is defined as a non-square scalar matrix for improved computational efficiency. Thus, the parameters $\mathbf{\Lambda}$ and $\mathbf{\Xi}$ are initialized as scalar matrices with all values of 1. Additionally, the diagonal matrix $\mathbf{\Gamma}$, with trainable diagonal entries, is initialized using a Gaussian distribution with a standard deviation of 1. To enhance convergence during the optimization process, a step decay variant for the dynamic learning rate is employed, systematically reducing the learning rate with a decay rate of 0.5 across all numerical experiments conducted in this research. Furthermore, during the sequential thresholding step, a tolerance value of $5\times 10^{-3}$ is utilized for $\epsilon$, effectively eliminating small $\mathbf{\Xi}$ and $\mathbf{\Lambda}$ values. A comprehensive overview of the hyperparameters for both the classical SINDy and ADAM-SINDy methods is provided in Table 1, detailing the parametric values chosen for the comparative analysis.
@@ 3 Numerical Experiments
This section details the benchmark problems used in this study, encompassing five systems of nonlinear ordinary differential equations (ODEs) and a sixth case involving a nonlinear partial differential equation (PDE). The final test case highlights the advantages of integrating trainable sparsifying hyperparameters $(\lambda)$ within the ADAM-SINDy framework. Each benchmark problem features a nonlinear parameter that poses challenges for the classical SINDy approach, and the SINDy-W method successfully identifies the governing dynamics for all cases as expected. These results underscore the challenges often encountered in practical SINDy applications (SINDy-W/o) when the correct nonlinear parameter is absent from the candidate library.
@@ 3.3 ABC Flow — Multiple Fourier Frequencies
In this section, we investigate the computation of multiple Fourier frequencies and amplitudes, focusing on the complex dynamics of an analytically defined chaotic fluid flow system. To facilitate this analysis, we examine the well-established benchmark problem known as the ABC (Arnold-Beltrami-Childress) flow, which is frequently employed to study intricate fluid dynamics and chaotic behavior in three-dimensional systems 54. The velocity components of the ABC flow, which adhere to incompressibility, are formulated as follows:
@@
Additionally, a supplementary experiment was conducted to explore the potential of utilizing the computational graph inherent to stochastic gradient descent (PyTorch environment) by employing $\mathbf{\Gamma}$ as a differential divergence operator, rather than as a scaling matrix for individual candidates. In fluid dynamics, ensuring a divergence-free velocity field, which aligns with the conservation of mass, is a critical constraint. By embedding this physical principle into the optimization process, ADAM-SINDy applies $\mathbf{\Gamma}$ as a divergence operator, thereby enforcing adherence to the continuity equation. This novel approach enables the identification of all six Fourier frequencies and their corresponding amplitudes within just 20 epochs, a substantial reduction in computational effort compared to the standard ADAM-SINDy method, which typically requires around 60,000 epochs when $\mathbf{\Gamma}$ is treated as a scaling matrix. This preliminary investigation suggests the potential for a physics-informed SINDy approach, which diverges from the primary focus of the current study and will be pursued in future work.
Figure 3: ABC flow example. (a) Pathline trajectory comparison showing both the ground truth and the SINDy-W/o method. (b) Loss plot illustrating the convergence behavior of the ADAM-SINDy method. A comparison of the governing equations is provided between the (c) ground truth and those identified by the (d) SINDy-W, (e) SINDy-W/o, and (f) ADAM-SINDy approaches. Optimized parameter values, highlighted in red, indicate results obtained through these methodologies. For the SINDy-W/o method, a total of 16 and 14 terms were identified for $\dot{x}$ and $\dot{z}$, reflecting the complexity of the derived dynamical system. In the ADAM-SINDy method, all optimized Fourier frequency values are approximated to 16 decimal places, consistently aligning with the value of $\pi$. Ground truth data is generated up to $t=20$ with the initial conditions $x_{0}=0.5$, $y_{0}=0.2$, and $z_{0}=1.0$, using a time step of 0.01.
@@ 3.4 Chemical Reaction Kinetics — Exponential Exponent
Next, the computation of an exponential exponent within a dynamical system is examined through a chemical reaction kinetics problem that employs an exponential approximation of the Arrhenius rate law, which characterizes the nonlinear relationship between reaction rates and temperature, particularly in the context of coupled first-order reactions 55. When the temperature increase due to an exothermic reaction is negligible compared to the ambient temperature, the reaction rate can be approximated as exponentially dependent on this incremental temperature change. By appropriately scaling the governing equations for mass and energy, the system is expressed as follows 55
@@
where $\alpha(t)$ represents the concentration of an intermediate reactant, $\theta(t)$ denotes the temperature rise, $k$ is the reaction rate constant, and $\mu$ corresponds to the reactant concentration. The model assumes that the states $\theta$ and $\alpha$ evolve on a faster time scale compared to the slowly changing reactant consumption $\mu$, invoking the pseudo-stationary-state hypothesis 56, thus $\mu$ interacting as a constant source term. This formulation offers insights into the temporal evolution of reactant concentrations and temperature, enabling the analysis of reaction kinetics and the impact of external perturbations on the reaction system.
Conventionally, the growth/decay rates $g$ and $h$ are assumed to be unity for the system under consideration 55, 56. However, this experiment adopts non-integer values for the parameters, specifically $g=1.015$ and $h=1.025$, while setting $k=0.07$ and $\mu=0.1$ to generate the ground truth data up to $t=200$ with initial conditions $\alpha_{0}=0$ and $\theta_{0}=5$ using a time step of $0.02$. Fig. 4 provides a comprehensive comparison of the classical SINDy-W, SINDy-W/o, and ADAM-SINDy methods in identifying the chemical reaction system. In this context, the SINDy-W/o method exhibits deviation from the ground truth trajectory at $\alpha$ close to zero, failing to accurately replicate the dynamics. In contrast, the ADAM-SINDy method successfully identifies the non-integer exponent values and other relevant coefficients perfectly aligning with the ground truth data, thereby demonstrating its efficacy in addressing systems characterized by exponential functions.

### v2

@@ 2.2 ADAM Optimization Augmented SINDy: ADAM-SINDy
$\mathcal{L}_{b}\leftarrow\left\|\dot{\mathbf{X}}_{b}-\bm{\Theta}(\mathbf{X}_{b};\bm{\Lambda})\bm{\Xi}\right\|_{2}^{2}+\lambda\left\|\bm{\Gamma}\bm{\Xi}\right\|_{1}$
       Update Parameters using Backpropagation:
$\bm{\Xi}_{n+1}\leftarrow\bm{\Xi}_{n}-\eta_{\bm{\Xi}}\nabla_{\bm{\Xi}}\mathcal{L}_{b}(\bm{\Theta}_{b},\bm{\Xi},\bm{\Lambda},\lambda,\bm{\Gamma})$
$\bm{\Lambda}_{n+1}\leftarrow\bm{\Lambda}_{n}-\eta_{\bm{\Lambda}}\nabla_{\bm{\Lambda}}\mathcal{L}_{b}(\bm{\Theta}_{b},\bm{\Xi},\bm{\Lambda},\lambda,\bm{\Gamma})$
        if Uniform Sparsity then
$\lambda_{n+1}\leftarrow\lambda_{n}+\eta_{\lambda}\nabla_{\lambda}\mathcal{L}_{b}(\bm{\Theta}_{b},\bm{\Xi},\bm{\Lambda},\lambda,\bm{\Gamma})$
        else
$\bm{\Gamma}_{n+1}\leftarrow\bm{\Gamma}_{n}+\eta_{\bm{\Gamma}}\nabla_{\bm{\Gamma}}\mathcal{L}_{b}(\bm{\Theta}_{b},\bm{\Xi},\bm{\Lambda},\lambda,\bm{\Gamma})$
       Thresholding: For each element $i$, set $\bm{\Xi}_{n+1}^{i},\bm{\Lambda}_{n+1}^{i}\leftarrow 0$ if $|\bm{\Xi}_{n+1}^{i}|,|\bm{\Lambda}_{n+1}^{i}|\leq\epsilon$
@@
 Optimized parameters $\bm{\Xi}^{*}$ and $\bm{\Lambda}^{*}$
@@ 2.3 Problem Statement
$\bm{\Theta}=\begin{bmatrix}1&\mathbf{X}^{\mathcal{A}}&\sin(\mathcal{B}\mathbf{X})&\cos(\mathcal{C}\mathbf{X})&\exp(\mathcal{D}\mathbf{X})&\mathbf{X}\otimes\sin(\mathcal{E}\mathbf{X})&\mathbf{X}\otimes\cos(\mathcal{F}\mathbf{X})&\mathbf{X}\otimes\exp(\mathcal{G}\mathbf{X})\end{bmatrix}\;,$ (9)
where $\mathbf{X}$ denotes the input data, and $\mathcal{A},\mathcal{B},\mathcal{C},\mathcal{D},\mathcal{E},\mathcal{F},\mathcal{G}$ represent the adaptive nonlinear parameters $(\bm{\Lambda})$ that are subject to optimization within the ADAM-SINDy framework. Here, $\mathcal{A}$ is the polynomial order, $\mathcal{B}$ and $\mathcal{C}$ represent Fourier frequencies, $\mathcal{D}$ corresponds to exponential growth/decay, $\mathcal{E}$ and $\mathcal{F}$ are the Fourier frequencies for Fourier terms combined with polynomial terms, and $\mathcal{G}$ represents the exponential growth/decay for the exponential terms combined with polynomial terms. Specifically, $\mathcal{A}$ is chosen to be 1, thereby encompassing polynomial terms up to the first order in the master library for all cases. However, in certain numerical experiments, $\mathcal{A}$ is rendered trainable when it corresponds to a non-integer rational exponent. It is noteworthy that the classical SINDy method relies on precise a priori knowledge of these parameters to successfully identify a sparse system.
@@
To investigate these limitations, two variants of the classical SINDy approach are explored: the SINDy-W method and the SINDy-W/o method. The SINDy-W method represents a fortunate scenario (unlikely in practical complex problems) where the precise value of the unknown nonlinear parameter is included within the range of candidate parameters $\bm{\Lambda}$. While such a scenario may hold in controlled or well-defined environments, they are often inadequate in real-world engineering applications, where the underlying parameter values may be unknown or difficult to estimate. Conversely, the SINDy-W/o method assumes no prior knowledge of the parametric values and applies a broad range of distributed parameter values to account for a wider spectrum of possibilities. In our benchmark problems, the range of parameters is selected such that the unknown parameter falls within the range. The SINDy-W/o approach is the more realistic practical version of SINDy, however, it often struggles in cases when the predefined parameter ranges fail to contain the “exact value” required to capture the system’s true behavior. This absence of prior knowledge frequently leads to less accurate, overly complex, and non-parsimonious models. Ultimately, these challenges underscore the need for flexibility and adaptability in parameter selection to ensure accurate system identification, motivating the use of the ADAM-SINDy framework, which dynamically adjusts parameter values during optimization and overcomes the limitations of the classical methods’ predefined parameter spaces.
@@
Here, we adopt the approach established in the classical SINDy method regarding the sparsity-promoting parameters $\lambda$ and $\bm{\Gamma}$  [34]. Specifically, $\bm{\Gamma}$ is treated as an identity matrix, with the $\lambda$ parameter being the sole variable used to control sparsity for the two variants considered: classical SINDy-W and SINDy-W/o methods. Notably, the numerical experiments conducted with the ADAM-SINDy framework indicate that the simultaneous optimization and utilization of both sparsity-promoting knobs is not beneficial.
@@
The ADAM-SINDy methodology is implemented using the PyTorch library, considering both batch-wise and full-batch optimization strategies, and employs the ADAM optimizer to facilitate the simultaneous optimization of all unknown variables involved in the sparse regression problem, including the parameters $\bm{\Lambda}$, the coefficients $\bm{\Xi}$, and the sparsity knob $\bm{\Gamma}$, with the parameter $\lambda$ fixed at a value of 1 and not subject to optimization. In this implementation, the parameter $\bm{\Lambda}\in\mathbb{R}^{m\times\hat{p}}$ is defined as a rectangular matrix associated with $\hat{p}$ nonlinear candidate parameters corresponding to each of the $m$ state variables. This structure facilitates vectorization, allowing the ADAM optimizer to be invoked only once for all nonlinear parameters rather than for each individual value. Consequently, while $\bm{\Lambda}$ parameters behave like scalar values algorithmically, it is implemented as a rectangular matrix in vectorized form for improved computational efficiency. Thus, the parameters $\bm{\Lambda}$ and $\bm{\Xi}$ are initialized as scalar matrices with all values of 1. Additionally, the diagonal matrix $\bm{\Gamma}$, with trainable diagonal entries, is initialized using a Gaussian distribution with a standard deviation of 1. To enhance convergence during the optimization process, a step decay variant for the dynamic learning rate is employed, systematically reducing the learning rate with a decay rate of 0.5 across all numerical experiments conducted in this research. Furthermore, during the sequential thresholding step, a tolerance value of $5\times 10^{-3}$ is utilized for $\epsilon$, effectively eliminating small $\bm{\Xi}$ and $\bm{\Lambda}$ values. A comprehensive overview of the hyperparameters for both the classical SINDy and ADAM-SINDy methods is provided in Table 1, detailing the parametric values chosen for the comparative analysis.
@@ 3 Numerical Experiments
This section details the benchmark problems used in this study, encompassing five systems of nonlinear ordinary differential equations (ODEs) and a sixth case involving a nonlinear partial differential equation (PDE). The final test case highlights the advantages of integrating trainable sparsifying hyperparameters, namely $\lambda$ and $\Gamma$, along with a sensitivity analysis. Each benchmark problem features a nonlinear parameter that poses challenges for the classical SINDy approach, and the SINDy-W method successfully identifies the governing dynamics for all cases as expected. These results underscore the challenges often encountered in practical SINDy applications (SINDy-W/o) when the correct nonlinear parameter is absent from the candidate library.
@@ 3.3 ABC Flow — Multiple Fourier Frequencies
In this section, we investigate the computation of multiple Fourier frequencies and amplitudes, focusing on the complex dynamics of an analytically defined chaotic fluid flow system. To facilitate this analysis, we examine the well-established benchmark problem known as the ABC (Arnold-Beltrami-Childress) flow, which is frequently employed to study intricate fluid dynamics and chaotic behavior in three-dimensional systems [59]. The velocity components of the ABC flow, which adhere to incompressibility, are formulated as follows:
@@
Additionally, a supplementary experiment was conducted to explore the potential of utilizing the computational graph inherent to stochastic gradient descent (PyTorch environment) to employ a differential divergence form in the $\ell_{1}$-norm regularization term. In fluid dynamics, ensuring a divergence-free velocity field, which aligns with the conservation of mass, is a critical constraint. To incorporate this physical principle into the optimization process, the sparse regression problem is redefined as follows
$\min_{\bm{\Xi}}\left[\|\dot{\mathbf{X}}-\bm{\Theta}(\mathbf{X}\textbf{;}\thinspace\bm{\Lambda})\thinspace\bm{\Xi}\|_{2}^{2}+\lambda\|\hat{\Gamma}\thinspace\bm{\Xi}\|_{1}\right]\;,$ (13)
where $\hat{\Gamma}$ represents the divergence equation, defined as $\partial\dot{x}_{m}/\partial x_{m}$ (where $\dot{x}_{m}=\bm{\Theta}(\mathbf{X};\bm{\Lambda}_{m})\bm{\Xi}_{m}$), computed efficiently using PyTorch’s autodifferentiation functionality. Notably, $\hat{\Gamma}$ is now a scalar value, distinct from the $\bm{\Gamma}$ matrix used previously. Consequently, the $\ell_{1}$-norm regularization term encapsulates both the divergence equation loss and sparsity ($\lambda>0$), thereby guiding the total loss function. This approach selectively retains candidate functions that satisfy the divergence-free condition, thereby ensuring compliance with the continuity equation within the ADAM-SINDy optimization framework. By fixing $\lambda$ to $1$, this novel approach enables the identification of all six Fourier frequencies and their corresponding amplitudes within just 40 epochs, as shown in Fig. 3, marking a substantial reduction in computational effort compared to the standard ADAM-SINDy method, which typically requires around 60,000 epochs when $\bm{\Gamma}$ is a weighting matrix. This preliminary investigation suggests the potential for a physics-informed SINDy approach, which diverges from the primary focus of the current study and will be pursued in future work.
Figure 3: ABC flow example. (a) Pathline trajectory comparison showing both the ground truth and the SINDy-W/o method. (b) Loss plot illustrating the convergence behavior of the ADAM-SINDy method with and without divergence-free condition. A comparison of the governing equations is provided between the (c) ground truth and those identified by the (d) SINDy-W, (e) SINDy-W/o, and (f) ADAM-SINDy approaches. Optimized parameter values, highlighted in red, indicate results obtained through these methodologies. For the SINDy-W/o method, a total of 16 and 14 terms were identified for $\dot{x}$ and $\dot{z}$, reflecting the complexity of the derived dynamical system. In the ADAM-SINDy method, all optimized Fourier frequency values are approximated to 16 decimal places, consistently aligning with the value of $\pi$. Ground truth data is generated up to $t=20$ with the initial conditions $x_{0}=0.5$, $y_{0}=0.2$, and $z_{0}=1.0$, using a time step of 0.01.
@@ 3.4 Chemical Reaction Kinetics — Exponential Exponent
Next, the computation of an exponential exponent within a dynamical system is examined through a chemical reaction kinetics problem that employs an exponential approximation of the Arrhenius rate law, which characterizes the nonlinear relationship between reaction rates and temperature, particularly in the context of coupled first-order reactions [60]. When the temperature increase due to an exothermic reaction is negligible compared to the ambient temperature, the reaction rate can be approximated as exponentially dependent on this incremental temperature change. By appropriately scaling the governing equations for mass and energy, the system is expressed as follows [60]
@@
where $\alpha(t)$ represents the concentration of an intermediate reactant, $\theta(t)$ denotes the temperature rise, $k$ is the reaction rate constant, and $\mu$ corresponds to the reactant concentration. The model assumes that the states $\theta$ and $\alpha$ evolve on a faster time scale compared to the slowly changing reactant consumption $\mu$, invoking the pseudo-stationary-state hypothesis [61], thus $\mu$ interacting as a constant source term. This formulation offers insights into the temporal evolution of reactant concentrations and temperature, enabling the analysis of reaction kinetics and the impact of external perturbations on the reaction system.
Conventionally, the growth/decay rates $g$ and $h$ are assumed to be unity for the system under consideration [60, 61]. However, this experiment adopts non-integer values for the parameters, specifically $g=1.015$ and $h=1.025$, while setting $k=0.07$ and $\mu=0.1$ to generate the ground truth data up to $t=200$ with initial conditions $\alpha_{0}=0$ and $\theta_{0}=5$ using a time step of $0.02$. Fig. 4 provides a comprehensive comparison of the classical SINDy-W, SINDy-W/o, and ADAM-SINDy methods in identifying the chemical reaction system. In this context, the SINDy-W/o method exhibits deviation from the ground truth trajectory at $\alpha$ close to zero, failing to accurately replicate the dynamics. In contrast, the ADAM-SINDy method successfully identifies the non-integer exponent values and other relevant coefficients perfectly aligning with the ground truth data, thereby demonstrating its efficacy in addressing systems characterized by exponential functions.